OpenStudy (anonymous):

dy/dx arcsin(1/x) my answer was -1 / x^2 sqrt(1 - 1/x^2) wolfram says this is correct, but my book shows answer as -1 / (x sqrt(x^2 - 1)) Im struggling to see how they are equivalent.

OpenStudy (anonymous):

i think it is just a matter of subtracting

OpenStudy (anonymous):

\[1-\frac{1}{x^2}\] \[\frac{x^2-1}{x^2}\]

OpenStudy (anonymous):

math books abhor complex fractions. it is clear now right?

OpenStudy (anonymous):

the \[x^2\] comes out of the square root as an x and cancels with one of them

OpenStudy (anonymous):

I think so, algebra is killing me here I think. Yes, that makes it clear thank you

OpenStudy (anonymous):

your last bit about what happens with x helped most. Thank you. I hate complex fractions. :(

OpenStudy (anonymous):

i can write it out if you like

OpenStudy (anonymous):

No, I see it now sqrt 1/x^2 is just 1/x x^2/x = x. got it.

OpenStudy (anonymous):

\[\frac{-1}{ x^2 \sqrt{1 - \frac{1}{x^2}}} \] \[\frac{-1}{x^2\sqrt{\frac{x^2-1}{x^2}}}\] \[\frac{-1}{\frac{x^2}{x}\sqrt{x^2-1}}\]

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

yw